Dublin University Rifle Club

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Mark Dennehy 167 Applewood Hts Greystones Co. Wicklow Ireland

Mark Dennehy, DURC Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000,  5/11/2001

A Chara,

Firstly, I would like to thank you for your patience with the running of the competition this year. Committee changes in the club meant that I had to learn how to do this on the run and many things did not go as planned. Notes have been taken: with luck these problems will be anticipated and prevented for the next competition.

Here (finally! ☺) are the results of the Trinity Top Ten for the year 2000-2001.

Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Score  Place
Andrew McFarland  EARC  24  24  20  25  25  19  26  23  19  26  231  1
Gary Duff  Comber  26  26  26  24  21  25  21  26  12  12  219  2
John Walsh  DRC  19  18  16  26  26  23  24  24  20  22  218  3
Robert Clarke  EARC  23  25  22  17  19  20  23  19  23  18  209  4
Richard Farrell  RRPC  20  14  25  23  16  21  22  22  21  25  209  4
Agnes Graham  Comber  15  16  17  21  20  26  18  25  26  24  208  6
Tom McKnight  EARC  17  23  18  18  13  15  20  13  25  13  175  7
Trevor McFarland  EARC  21  19  9  20  15  12  16  15  24  20  171  8
Ronan Power  RRPC  11  13  12  13  24  16  25  17  16  16  163  9
Elizabeth Rodgers  EARC  12  8  10  14  10  22  15  21  22  23  157  10
Raymond Reilly  DURC  22  21  24  23  22  0  9  8  10  10  149  11
Ivor Fletcher  EARC  9  10  13  15  7  17  13  18  17  17  136  12
Maruice Davis  EARC  10  5  21  10  18  10  19  12  14  15  134  13
Dave Cullen  RRPC  25  20  23  19  23  18  1  1  1  1  132  14
Robin Crockett Jr.  EARC  5  22  0  11  12  14  11  20  15  19  129  15
Jim Dace  EARC  18  9  6  9  17  13  14  14  13  14  127  16
John Keeney  DURC  7  11  11  12  8  8  12  16  18  21  124  17
Mark Dennehy  DURC  4  12  14  16  11  11  17  11  9  9  114  18
Kealan Symes  RRPC  16  17  19  7  14  24  3  2  3  3  108  19
Dave Trumper  RRPC  8  4  7  8  9  9  2  10  2  2  61  20
John McVey  EARC  14  15  8  0  0  0  6  5  6  6  60  21
Nick West  RRPC  13  6  15  0  0  0  4  3  4  4  49  22
Willie O Connor  DURC  0  0  0  0  0  0  10  9  11  11  41  23
Alan Smyth  DURC  6  7  0  0  0  0  7  6  7  7  40  24
Ian Marnane  DURC  0  0  0  0  0  0  8  7  8  8  31  25
Dave Cooney  FSC  0  0  0  0  0  0  5  4  5  5  19  26

Once again, many thanks for your patience, and I hope that we can get as many entrants or more for the next Top Ten, which will be marking our 40^{th} anniversary year. Entry forms for this will be sent out shortly.

Mise le meas,

Mark Dennehy, DURC Secretary

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Dave Cooney Ardnaglass Corrig Road Dun Laoghaire

Mark Dennehy, DURC Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000,  5/11/2001

A Chara,

Firstly, I would like to thank you for your patience with the running of the competition this year. Committee changes in the club meant that I had to learn how to do this on the run and many things did not go as planned. Notes have been taken: with luck these problems will be anticipated and prevented for the next competition.

Here (finally! ☺) are the results of the Trinity Top Ten for the year 2000-2001.

Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Score  Place
Andrew McFarland  EARC  24  24  20  25  25  19  26  23  19  26  231  1
Gary Duff  Comber  26  26  26  24  21  25  21  26  12  12  219  2
John Walsh  DRC  19  18  16  26  26  23  24  24  20  22  218  3
Robert Clarke  EARC  23  25  22  17  19  20  23  19  23  18  209  4
Richard Farrell  RRPC  20  14  25  23  16  21  22  22  21  25  209  4
Agnes Graham  Comber  15  16  17  21  20  26  18  25  26  24  208  6
Tom McKnight  EARC  17  23  18  18  13  15  20  13  25  13  175  7
Trevor McFarland  EARC  21  19  9  20  15  12  16  15  24  20  171  8
Ronan Power  RRPC  11  13  12  13  24  16  25  17  16  16  163  9
Elizabeth Rodgers  EARC  12  8  10  14  10  22  15  21  22  23  157  10
Raymond Reilly  DURC  22  21  24  23  22  0  9  8  10  10  149  11
Ivor Fletcher  EARC  9  10  13  15  7  17  13  18  17  17  136  12
Maruice Davis  EARC  10  5  21  10  18  10  19  12  14  15  134  13
Dave Cullen  RRPC  25  20  23  19  23  18  1  1  1  1  132  14
Robin Crockett Jr.  EARC  5  22  0  11  12  14  11  20  15  19  129  15
Jim Dace  EARC  18  9  6  9  17  13  14  14  13  14  127  16
John Keeney  DURC  7  11  11  12  8  8  12  16  18  21  124  17
Mark Dennehy  DURC  4  12  14  16  11  11  17  11  9  9  114  18
Kealan Symes  RRPC  16  17  19  7  14  24  3  2  3  3  108  19
Dave Trumper  RRPC  8  4  7  8  9  9  2  10  2  2  61  20
John McVey  EARC  14  15  8  0  0  0  6  5  6  6  60  21
Nick West  RRPC  13  6  15  0  0  0  4  3  4  4  49  22
Willie O Connor  DURC  0  0  0  0  0  0  10  9  11  11  41  23
Alan Smyth  DURC  6  7  0  0  0  0  7  6  7  7  40  24
Ian Marnane  DURC  0  0  0  0  0  0  8  7  8  8  31  25
Dave Cooney  FSC  0  0  0  0  0  0  5  4  5  5  19  26

Mise le meas,

Mark Dennehy, DURC Secretary

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

John K. Walsh Carjon Lodge Cross Chapel Blessington Co. Wicklow

Mark Dennehy, DURC Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000,  5/11/2001

A Chara,

Firstly, I would like to thank you for your patience with the running of the competition this year. Committee changes in the club meant that I had to learn how to do this on the run and many things did not go as planned. Notes have been taken: with luck these problems will be anticipated and prevented for the next competition.

Here (finally! ☺) are the results of the Trinity Top Ten for the year 2000-2001.

Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Score  Place
Andrew McFarland  EARC  24  24  20  25  25  19  26  23  19  26  231  1
Gary Duff  Comber  26  26  26  24  21  25  21  26  12  12  219  2
John Walsh  DRC  19  18  16  26  26  23  24  24  20  22  218  3
Robert Clarke  EARC  23  25  22  17  19  20  23  19  23  18  209  4
Richard Farrell  RRPC  20  14  25  23  16  21  22  22  21  25  209  4
Agnes Graham  Comber  15  16  17  21  20  26  18  25  26  24  208  6
Tom McKnight  EARC  17  23  18  18  13  15  20  13  25  13  175  7
Trevor McFarland  EARC  21  19  9  20  15  12  16  15  24  20  171  8
Ronan Power  RRPC  11  13  12  13  24  16  25  17  16  16  163  9
Elizabeth Rodgers  EARC  12  8  10  14  10  22  15  21  22  23  157  10
Raymond Reilly  DURC  22  21  24  23  22  0  9  8  10  10  149  11
Ivor Fletcher  EARC  9  10  13  15  7  17  13  18  17  17  136  12
Maruice Davis  EARC  10  5  21  10  18  10  19  12  14  15  134  13
Dave Cullen  RRPC  25  20  23  19  23  18  1  1  1  1  132  14
Robin Crockett Jr.  EARC  5  22  0  11  12  14  11  20  15  19  129  15
Jim Dace  EARC  18  9  6  9  17  13  14  14  13  14  127  16
John Keeney  DURC  7  11  11  12  8  8  12  16  18  21  124  17
Mark Dennehy  DURC  4  12  14  16  11  11  17  11  9  9  114  18
Kealan Symes  RRPC  16  17  19  7  14  24  3  2  3  3  108  19
Dave Trumper  RRPC  8  4  7  8  9  9  2  10  2  2  61  20
John McVey  EARC  14  15  8  0  0  0  6  5  6  6  60  21
Nick West  RRPC  13  6  15  0  0  0  4  3  4  4  49  22
Willie O Connor  DURC  0  0  0  0  0  0  10  9  11  11  41  23
Alan Smyth  DURC  6  7  0  0  0  0  7  6  7  7  40  24
Ian Marnane  DURC  0  0  0  0  0  0  8  7  8  8  31  25
Dave Cooney  FSC  0  0  0  0  0  0  5  4  5  5  19  26

Mise le meas,

Mark Dennehy, DURC Secretary

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Agnes Graham 114 Towerview Crescent Bangor Co. Down BT19 6A2

Mark Dennehy, DURC Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000,  5/11/2001

A Chara,

Firstly, I would like to thank you for your patience with the running of the competition this year. Committee changes in the club meant that I had to learn how to do this on the run and many things did not go as planned. Notes have been taken: with luck these problems will be anticipated and prevented for the next competition.

Here (finally! ☺) are the results of the Trinity Top Ten for the year 2000-2001.

Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Score  Place
Andrew McFarland  EARC  24  24  20  25  25  19  26  23  19  26  231  1
Gary Duff  Comber  26  26  26  24  21  25  21  26  12  12  219  2
John Walsh  DRC  19  18  16  26  26  23  24  24  20  22  218  3
Robert Clarke  EARC  23  25  22  17  19  20  23  19  23  18  209  4
Richard Farrell  RRPC  20  14  25  23  16  21  22  22  21  25  209  4
Agnes Graham  Comber  15  16  17  21  20  26  18  25  26  24  208  6
Tom McKnight  EARC  17  23  18  18  13  15  20  13  25  13  175  7
Trevor McFarland  EARC  21  19  9  20  15  12  16  15  24  20  171  8
Ronan Power  RRPC  11  13  12  13  24  16  25  17  16  16  163  9
Elizabeth Rodgers  EARC  12  8  10  14  10  22  15  21  22  23  157  10
Raymond Reilly  DURC  22  21  24  23  22  0  9  8  10  10  149  11
Ivor Fletcher  EARC  9  10  13  15  7  17  13  18  17  17  136  12
Maruice Davis  EARC  10  5  21  10  18  10  19  12  14  15  134  13
Dave Cullen  RRPC  25  20  23  19  23  18  1  1  1  1  132  14
Robin Crockett Jr.  EARC  5  22  0  11  12  14  11  20  15  19  129  15
Jim Dace  EARC  18  9  6  9  17  13  14  14  13  14  127  16
John Keeney  DURC  7  11  11  12  8  8  12  16  18  21  124  17
Mark Dennehy  DURC  4  12  14  16  11  11  17  11  9  9  114  18
Kealan Symes  RRPC  16  17  19  7  14  24  3  2  3  3  108  19
Dave Trumper  RRPC  8  4  7  8  9  9  2  10  2  2  61  20
John McVey  EARC  14  15  8  0  0  0  6  5  6  6  60  21
Nick West  RRPC  13  6  15  0  0  0  4  3  4  4  49  22
Willie O Connor  DURC  0  0  0  0  0  0  10  9  11  11  41  23
Alan Smyth  DURC  6  7  0  0  0  0  7  6  7  7  40  24
Ian Marnane  DURC  0  0  0  0  0  0  8  7  8  8  31  25
Dave Cooney  FSC  0  0  0  0  0  0  5  4  5  5  19  26

Mise le meas,

Mark Dennehy, DURC Secretary

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Ivor Fletcher 127 Belfast Road Antrim Co. Antrim BT41 2AZ

Mark Dennehy, DURC Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000,  5/11/2001

A Chara,

Firstly, I would like to thank you for your patience with the running of the competition this year. Committee changes in the club meant that I had to learn how to do this on the run and many things did not go as planned. Notes have been taken: with luck these problems will be anticipated and prevented for the next competition.

Here (finally! ☺) are the results of the Trinity Top Ten for the year 2000-2001.

Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Score  Place
Andrew McFarland  EARC  24  24  20  25  25  19  26  23  19  26  231  1
Gary Duff  Comber  26  26  26  24  21  25  21  26  12  12  219  2
John Walsh  DRC  19  18  16  26  26  23  24  24  20  22  218  3
Robert Clarke  EARC  23  25  22  17  19  20  23  19  23  18  209  4
Richard Farrell  RRPC  20  14  25  23  16  21  22  22  21  25  209  4
Agnes Graham  Comber  15  16  17  21  20  26  18  25  26  24  208  6
Tom McKnight  EARC  17  23  18  18  13  15  20  13  25  13  175  7
Trevor McFarland  EARC  21  19  9  20  15  12  16  15  24  20  171  8
Ronan Power  RRPC  11  13  12  13  24  16  25  17  16  16  163  9
Elizabeth Rodgers  EARC  12  8  10  14  10  22  15  21  22  23  157  10
Raymond Reilly  DURC  22  21  24  23  22  0  9  8  10  10  149  11
Ivor Fletcher  EARC  9  10  13  15  7  17  13  18  17  17  136  12
Maruice Davis  EARC  10  5  21  10  18  10  19  12  14  15  134  13
Dave Cullen  RRPC  25  20  23  19  23  18  1  1  1  1  132  14
Robin Crockett Jr.  EARC  5  22  0  11  12  14  11  20  15  19  129  15
Jim Dace  EARC  18  9  6  9  17  13  14  14  13  14  127  16
John Keeney  DURC  7  11  11  12  8  8  12  16  18  21  124  17
Mark Dennehy  DURC  4  12  14  16  11  11  17  11  9  9  114  18
Kealan Symes  RRPC  16  17  19  7  14  24  3  2  3  3  108  19
Dave Trumper  RRPC  8  4  7  8  9  9  2  10  2  2  61  20
John McVey  EARC  14  15  8  0  0  0  6  5  6  6  60  21
Nick West  RRPC  13  6  15  0  0  0  4  3  4  4  49  22
Willie O Connor  DURC  0  0  0  0  0  0  10  9  11  11  41  23
Alan Smyth  DURC  6  7  0  0  0  0  7  6  7  7  40  24
Ian Marnane  DURC  0  0  0  0  0  0  8  7  8  8  31  25
Dave Cooney  FSC  0  0  0  0  0  0  5  4  5  5  19  26

Mise le meas,

Mark Dennehy, DURC Secretary

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Karen Farrell 41 Deerpark Walk Blessington Co. Wicklow

Mark Dennehy, DURC Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000,  5/11/2001

A Chara,

Firstly, I would like to thank you for your patience with the running of the competition this year. Committee changes in the club meant that I had to learn how to do this on the run and many things did not go as planned. Notes have been taken: with luck these problems will be anticipated and prevented for the next competition.

Here (finally! ☺) are the results of the Trinity Top Ten for the year 2000-2001.

Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Score  Place
Andrew McFarland  EARC  24  24  20  25  25  19  26  23  19  26  231  1
Gary Duff  Comber  26  26  26  24  21  25  21  26  12  12  219  2
John Walsh  DRC  19  18  16  26  26  23  24  24  20  22  218  3
Robert Clarke  EARC  23  25  22  17  19  20  23  19  23  18  209  4
Richard Farrell  RRPC  20  14  25  23  16  21  22  22  21  25  209  4
Agnes Graham  Comber  15  16  17  21  20  26  18  25  26  24  208  6
Tom McKnight  EARC  17  23  18  18  13  15  20  13  25  13  175  7
Trevor McFarland  EARC  21  19  9  20  15  12  16  15  24  20  171  8
Ronan Power  RRPC  11  13  12  13  24  16  25  17  16  16  163  9
Elizabeth Rodgers  EARC  12  8  10  14  10  22  15  21  22  23  157  10
Raymond Reilly  DURC  22  21  24  23  22  0  9  8  10  10  149  11
Ivor Fletcher  EARC  9  10  13  15  7  17  13  18  17  17  136  12
Maruice Davis  EARC  10  5  21  10  18  10  19  12  14  15  134  13
Dave Cullen  RRPC  25  20  23  19  23  18  1  1  1  1  132  14
Robin Crockett Jr.  EARC  5  22  0  11  12  14  11  20  15  19  129  15
Jim Dace  EARC  18  9  6  9  17  13  14  14  13  14  127  16
John Keeney  DURC  7  11  11  12  8  8  12  16  18  21  124  17
Mark Dennehy  DURC  4  12  14  16  11  11  17  11  9  9  114  18
Kealan Symes  RRPC  16  17  19  7  14  24  3  2  3  3  108  19
Dave Trumper  RRPC  8  4  7  8  9  9  2  10  2  2  61  20
John McVey  EARC  14  15  8  0  0  0  6  5  6  6  60  21
Nick West  RRPC  13  6  15  0  0  0  4  3  4  4  49  22
Willie O Connor  DURC  0  0  0  0  0  0  10  9  11  11  41  23
Alan Smyth  DURC  6  7  0  0  0  0  7  6  7  7  40  24
Ian Marnane  DURC  0  0  0  0  0  0  8  7  8  8  31  25
Dave Cooney  FSC  0  0  0  0  0  0  5  4  5  5  19  26

Mise le meas,

Mark Dennehy, DURC Secretary